OpenStudy (anonymous):

A 30-foot ladder leans against a wall at a 77° angle of elevation. Find the distance from the ladder's base to the wall. Can I use the Pythagorean Theorem to solve this problem?

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

pythagorean can only be done when you're given two sides

OpenStudy (anonymous):

How do I solve it?

OpenStudy (lgbasallote):

draw this thingy first  does that give you an idea how to solve for x

OpenStudy (anonymous):

no idea

OpenStudy (lgbasallote):

hint \[\cos \theta = \frac{adjacent}{hypotenuse}\] \[\sin \theta = \frac{opposite}{hypotenuse}\] \[\tan \theta = \frac{opposite}{adjacent}\] any clue now?

OpenStudy (anonymous):

LOL I may sound dumb but I don't understand

OpenStudy (lgbasallote):

hmm...you know what hypotenuse is right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Ok this is my question you are telling me that I can't use the pythagorean Theorem but that's what the book is showing me.

OpenStudy (lgbasallote):

you can only use pythagorean theorem when you're given TWO sides and you're looking for the THIRD side

OpenStudy (anonymous):

I'm tired, my brain is tire. Ok can you go step by step

OpenStudy (lgbasallote):

but here you're given ONE side and ONE angle

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

anyway...the hypotenuse is the longest side. it is the side opposite the right angle...so in this case what's the hypotenuse?

OpenStudy (anonymous):

30 FT